OpenStudy (anonymous):

Expand the following function using Taylor Series about a = pi / 4 \[\cos 2x\]

OpenStudy (anonymous):

cos 2x comes where?

OpenStudy (anonymous):

f(x) = cos 2x. expand using taylor series about a = pi / 4

OpenStudy (anonymous):

cos(x) = 1/√2 - 1/√2(x-π/4) - 1/√2(x-π/4)^2/2 + 1/√2(x-π/4)^3/6 + 1/√2(x-π/4)^4/24 + .

OpenStudy (anonymous):

can you use the equation thing. i cant make up the answer.

OpenStudy (amistre64):

i know a cos(x) but im not sure where the 2x comes into play

OpenStudy (amistre64):

cos 1 - sin 0 - cos 1 would suggest 1 -(x-pi/4)^2 ... even powers to me alternating plus minuses divided by the exponents factorial but then again there is that "2" to place

OpenStudy (anonymous):

hint: evaluate the derivatives

OpenStudy (amistre64):

hmm, so use 2x in the derivatives ...

OpenStudy (amistre64):

cos (2x) -2sin (2x) -4cos(2x) -6 cos(2x) like that

OpenStudy (amistre64):

2(pi/4) = pi/2 so all these are the 90 degrees

OpenStudy (amistre64):

which makes sin the dominant one if im reading it right

OpenStudy (anonymous):

yes youre getting there

OpenStudy (amistre64):

if I were to take a gander, and chk it with the wolf: \[cos(2x)~-\frac{2}{1!}(x-\frac{pi}{4})+\frac{6}{3!}(x-\frac{pi}{4})^3-\frac{10}{5!}(x-\frac{pi}{4})^5+\frac{14}{7!}(x-\frac{pi}{4})^7 ...\]

OpenStudy (amistre64):

it ate the = sign :)

OpenStudy (anonymous):

Theres something wrong but its nearly done

OpenStudy (amistre64):

\[\sum_{n=1}^{\infty} (-1)^{n}\frac{4n-2}{(2n-1)!}(x-\frac{pi}{4})^{2n-1}\] maybe :) but thats all i got

OpenStudy (anonymous):

Its actually:\[\sum_{n = 0}^{\infty}\frac{(-1)^{n}2^{2n + 1}(x - \frac{\pi}{4})^{2n + 1}}{(2n + 1)!}\]

OpenStudy (anonymous):

or\[\cos 2x = -\frac{2}{1!}(x - \frac{\pi}{4}) + \frac{2^{3}}{3!}(x - \frac{\pi}{4})^{3} - ... \]

OpenStudy (amistre64):

2*2*2*2 ..... yeah, sooo close :)

OpenStudy (anonymous):

you might want to try the puzzle i posted. Not related to calculus though.

OpenStudy (anonymous):

find Taylor expansion of X-1/X+1 , at Xo = 0